# Unsupervised Post-Training for Multi-Modal LLM Reasoning via GRPO

### v1

@@ 4.4 Ablation Study
To evaluate the generality and effectiveness of MM-UPT, we conduct an ablation study across a range of backbone models beyond the primary Qwen2.5-VL-7B [1]. Specifically, we apply MM-UPT to several state-of-the-art models of varying scales, including Qwen2.5-VL-3B [1], MM-Eureka-7B [28], and ThinkLite-VL-7B [44]. All models are post-trained using MM-UPT on the Geometry3K dataset [23], without access to any labels. As summarized in Table 3, MM-UPT consistently improves the performance of all tested models on average, despite the absence of supervision during post-training. Notably, ThinkLite-VL-7B with MM-UPT achieves the highest average score (54.07), and shows substantial gains on the MathVista [24] benchmark, reaching a score of 74.70. In addition, Qwen2.5-VL-3B, the smallest model in our study, also benefits well from MM-UPT (+7.4% on average), demonstrating the robustness and adaptability of MM-UPT for performance enhancement. These results collectively reveal that MM-UPT can be easily applied to various multi-modal models to enable consistent self-improvement.
@@
Qwen2.5-VL-7B + MM-UPT 27.33 42.46 68.50 66.61 51.23 (3.6$\uparrow$)
MM-Eureka-7B 28.06 50.46 69.40 64.48 53.10
MM-Eureka-7B + MM-UPT 28.95 50.63 69.10 66.44 53.78 (1.3$\uparrow$)
@@ 5 Deeper Analysis
Going beyond standard benchmarking, we conduct a deeper analysis to investigate MM-UPT’s training dynamics (Section 5.1) and performance boundaries (Section 5.2 and Section 5.3). This helps better understand its behavior and potential applications.
5.1 Training Dynamics
$p(c_{j}|q,I)=\frac{|c_{j}|}{G},$
$H=-\sum_{c_{j}\in\{C\}}p(c_{j}|q,I)\log p(c_{j}|q,I).$
@@ 5.3 When Might MM-UPT Fail?
According to the analysis in Section 5.2, it reveals that the effectiveness of MM-UPT diminishes when the model lacks sufficient prior knowledge of the target dataset. To show that, we apply MM-UPT to ThinkLite-11K [44] dataset using Qwen2.5-VL-7B [1]. ThinkLite-11K is collected via difficulty-aware sampling that only retains samples that the model rarely answers correctly. Thus, this setting reflects a scenario where the model is more likely to be wrong than right. In such cases, majority voting amplifies incorrect answers rather than filtering them, leading to degraded performance. As shown in Table 4, applying MM-UPT to ThinkLite-11K results in a significant drop in accuracy across all benchmarks. This suggests that majority voting fails to provide reliable reward signals when the model has limited prior understanding of the domain. To address this issue, alternative forms of algorithms in self-improvement [49, 26, 7, 56] using more fine-grained and complex self-rewarding methods, such as LLM-as-Judge [65] and model collaboration [20], may be necessary. Note that our work represents an initial attempt at self-improvement in MLLMs via GRPO. We believe that these algorithms are complementary to our approach and could be integrated into our framework, which we leave as future work.
@@ 6 Conclusion
In this paper, we introduce MM-UPT (Multi-Modal Unsupervised Post-Training), a framework that enables MLLMs to self-improve using unlabeled multi-modal data without any external supervision. By leveraging majority voting as a reward mechanism within the GRPO algorithm, our method encourages models to converge toward more consistent responses across multi-modal reasoning tasks. Our experiments demonstrate that MM-UPT significantly improves the performance of various MLLMs across several multi-modal reasoning benchmarks without requiring human annotated labels or external reward models. Generally, MM-UPT is best viewed as a post-training refinement strategy rather than a substitute for supervised training. We recommend first training the model using labeled data to ensure it achieves basic competence on various tasks. After that, MM-UPT can be used to further exploit the model’s internal knowledge by reinforcing consistent predictions. Future work could explore other fine-grained methods to provide pseudo-reward signals based on our framework, and investigate the scaling laws of unsupervised post-training using synthetic data.
Acknowledgement
@@ A.3 Standard Training Datasets
Appendix B Compute Resources

### v2

@@ 4.4 Ablation Study
To evaluate the generality and effectiveness of MM-UPT, we conduct an ablation study across a range of backbone models beyond the primary Qwen2.5-VL-7B [1]. Specifically, we apply MM-UPT to several state-of-the-art models of varying scales, including Qwen2.5-VL-3B [1], MM-Eureka-7B [31], and ThinkLite-VL-7B [48]. All models are post-trained using MM-UPT on the Geometry3K dataset [25], without access to any labels (i.e., Scenario 1). As summarized in Table 3, MM-UPT consistently improves the performance of all tested models on average, despite the absence of supervision during post-training. Notably, ThinkLite-VL-7B with MM-UPT achieves the highest average score (54.07), and shows substantial gains on the MathVista [27] benchmark, reaching a score of 74.70. In addition, Qwen2.5-VL-3B, the smallest model in our study, also benefits well from MM-UPT (+7.4% on average), demonstrating the robustness and adaptability of MM-UPT for performance enhancement. These results collectively reveal that MM-UPT can be easily applied to various multi-modal models to enable consistent self-improvement.
Moreover, our results also show that MM-UPT is compatible with supervised GRPO. For instance, MM-Eureka-7B was already tuned with supervised GRPO on the K12 dataset, yet applying MM-UPT on a new unlabeled dataset (Geometry3K) further improved its average score from 53.10 to 53.78 (Table 3). Similarly, ThinkLite-VL-7B was trained with supervised GRPO on the ThinkLite dataset, and applying MM-UPT again boosted its average score to 54.07, including a substantial gain on MathVista (74.70). These results highlight the practical value of MM-UPT as a lightweight refinement step that remains effective even for models already optimized with supervised GRPO, enabling them to leverage new unlabeled data for further improvement.
@@
[…]
$M$-Eureka-7B 28.06 50.46 69.40 64.48 53.10
[…]
@@ 5 Deeper Analysis
Going beyond standard benchmarking, we conduct a deeper analysis to investigate MM-UPT’s performance boundaries (Section 5.1 and Section 5.2) and tradeoffs (Section 5.3). This helps better understand its behavior and potential applications.
@@ 5.2 When Might MM-UPT Fail?
According to the analysis in Section 5.1, it reveals that the effectiveness of MM-UPT diminishes when the model lacks sufficient prior knowledge of the target dataset. To show that, we apply MM-UPT to ThinkLite-11K [48] dataset using Qwen2.5-VL-7B [1]. ThinkLite-11K is collected via difficulty-aware sampling that only retains samples that the model rarely answers correctly. Thus, this setting reflects a scenario where the model is more likely to be wrong than right. In such cases, majority voting amplifies incorrect answers rather than filtering them, leading to degraded performance. As shown in Table 4, applying MM-UPT to ThinkLite-11K results in a significant drop in accuracy across all benchmarks. This suggests that majority voting fails to provide reliable reward signals when the model has limited prior understanding of the domain. To address this issue, alternative forms of algorithms using more fine-grained and complex rewarding methods, such as LLM-as-Judge [76, 54] and model collaboration [21, 8], may be necessary. Note that our work represents an initial attempt at self-improvement in MLLMs via GRPO, and we believe that these algorithms are complementary to our approach and could be integrated into our framework in the future.
5.3 Trade-Offs in MM-UPT
Although MM-UPT improves pass@1 accuracy across multiple benchmarks, we also observe a consistent decrease in pass@n (for large n, e.g., pass@10) performance in Table 5. This reflects a common trade-off between accuracy and diversity in reinforcement learning for multi-modal models. Similar findings have been reported in supervised GRPO [66] (also shown in Table 5), where models tend to collapse onto high-consensus reasoning patterns, thereby reducing output diversity. In MM-UPT, this effect is further amplified because the implicit reward signal from majority voting encourages the model to favor dominant responses, which may suppress minority answers that are occasionally correct. That is said, MM-UPT let the model reinforce problems it can already solve, thereby forgoing the opportunity to tackle more challenging ones and abandoning exploration. While pass@1 is typically the most relevant metric in real-world applications, mitigating the decline of pass@n remains an important open problem for future work.
Table 5: Comparison of pass@10 across different benchmarks. Results show that MM-UPT improves single-response accuracy (pass@1, Table 1) but reduces response diversity, leading to a drop in pass@10. Supervised GRPO alleviates this issue to some extent.
Model MathVision MathVerse MathVista We-Math Avg
Qwen2.5-VL-7B 0.6556 0.7307 0.8730 0.9420 0.8003
Qwen2.5-VL-7B + MM-UPT 0.5661 0.6477 0.8240 0.8621 0.7250
Qwen2.5-VL-7B + Supervised GRPO 0.6164 0.6726 0.8570 0.9075 0.7634
Another key consideration is the trade-off between training and inference costs. A straightforward alternative to MM-UPT is to apply majority voting directly at inference time, which can also boost accuracy by aggregating multiple outputs. However, inference-time ensembling incurs substantial computational overhead, as $n$ samples must be generated per query, making it expensive and often impractical at scale. By contrast, MM-UPT shifts this cost into a one-time training stage: after refinement, the model produces stronger single-pass outputs, avoiding the need for repeated sampling during deployment. This distinction highlights different usage scenarios: MM-UPT is particularly beneficial when inference efficiency and scalability are critical, whereas inference-time ensembling may be preferable when training resources are limited.
In summary, MM-UPT offers a simple unsupervised post-training framework but also entails clear trade-offs. Balancing accuracy and diversity, and choosing between training-time and inference-time costs, are crucial aspects to consider and represent promising directions for future exploration.
@@ 6 Conclusion
In this work, we formalize a third-stage post-training paradigm for multi-modal large language models after SFT and RL, termed Unsupervised Post-Training (UPT). We instantiate this paradigm through MM-UPT, a simple yet effective framework that leverages majority voting as a self-rewarding mechanism within the GRPO algorithm, guiding models toward consistent and high-confidence responses on multi-modal reasoning tasks. This can be used to further exploit the model’s internal knowledge by reinforcing consistent predictions, and thus enables self-improvement without any external supervision. Extensive experiments across multiple benchmarks demonstrate that this method effectively enhances the reasoning performance of strong MLLMs without relying on labeled data or external reward models. Furthermore, we extend the framework to a data self-generation setting, showing that synthetic questions produced by the model itself can further boost performance, revealing a scalable path toward autonomous self-improvement. Future work could explore other fine-grained methods to provide pseudo-reward signals based on our framework, and investigate the scaling laws of unsupervised post-training using synthetic data.
Acknowledgements and Disclosure of Funding
@@ A.3 Standard Training Datasets
Appendix B Additional Experiments
B.1 Training Dynamics
$p(c_{j}|q,I)=\frac{|c_{j}|}{G},$
$H=-\sum_{c_{j}\in\{C\}}p(c_{j}|q,I)\log p(c_{j}|q,I).$
B.2 Generalization Beyond Multimodal Mathematical Reasoning
A potential concern is that by encouraging convergence toward high-consensus answers, MM-UPT might reduce response diversity and overspecialize on mathematical reasoning, potentially harming performance on other tasks. To investigate whether this method suffers from a negative impact on broader generalization, we extend our evaluation to two non-mathematical visual question answering benchmarks: ChartQA [30], which tests visual perception over charts, and IconQA [26], which focuses on abstract diagram understanding. In particular, we evaluate the Qwen2.5-VL-7B model trained with MM-UPT on the MMR1 dataset. As shown in Table 6, our method not only avoids performance degradation but leads to notable improvements on both benchmarks. This suggests that the underlying mechanism of MM-UPT—reinforcing the generation of more reliable and consistent answers—is a beneficial trait that positively transfers to other domains. The results indicate that the induced consistency does not degrade, and can even enhance, the model’s general utility on related visual understanding tasks.
Table 6: Performance on non-mathematical VQA benchmarks. We evaluate Qwen2.5-VL-7B before and after applying MM-UPT on the MMR1 dataset. Scores are reported as accuracy.
Models ChartQA IconQA
Qwen2.5-VL-7B 71.96 54.20
[…]
B.3 Adaptability to Language Tasks
Furthermore, to assess the adaptability of our MM-UPT framework to purely language-based domains, we apply MM-UPT to a language model, Qwen2.5-MATH-7B [58], trained on the DAPO-17K dataset [63] using the same self-rewarding mechanism. We evaluate the resulting model on two popular pure-math benchmarks: MATH [14] and Omni-MATH [9]. As shown in Table 7, MM-UPT leads to substantial improvements on both benchmarks. These results strongly suggest that our framework is not limited to multi-modal settings and can be effectively extended to purely language-based reasoning tasks, provided the base model demonstrates sufficient initial competency.
Table 7: Performance on pure-language mathematical reasoning benchmarks. We evaluate Qwen2.5-MATH-7B before and after applying our unsupervised post-training method.
Models MATH Omni-MATH
Qwen2.5-MATH-7B 51.20 18.09
[…]
Appendix C Compute Resources